OpenStudy (anonymous):

Let f(x)=10x-10. Find the value of (f o f^-1)(-10)

OpenStudy (anonymous):

f(x) is a one-to-one function (it's a line) so (f o f^-1)(-10) = -10

OpenStudy (anonymous):

basically, inverses "undo" each other, so you end up with the same value

OpenStudy (anonymous):

oh ok and so i plug the -10 into the original equation or would that be the answer?

OpenStudy (jhannybean):

Nope.\[f(f^{-1}(-10)) = x\]\[-10=x\] That's basically what is going on. \[f(-10)=-10x-10 = -110 \ne -10 \]

OpenStudy (jhannybean):

Sorry, forgot to substitute by -10 in for x. \(f(-10)=-10(-10)-10 = -110\)*

OpenStudy (unklerhaukus):

\[f(x) = 10x-10\\[3ex] x=10\Big(f^{-1}(x)\Big)-10\\ x+10 = 10(f^{-1}(x))\\ f^{-1}(x)=\frac {x+10}{10}\\[3ex]f\circ f^{-1}(x)=f\Big(\frac {x+10}{10}\Big)\\ \qquad\qquad\quad = 10\Big(\frac {x+10}{10}\Big)-10\\ \qquad\qquad\quad = x+10-10\\ \qquad\qquad\quad =x\\[2ex] f\circ f^{-1}(-10) =-10\]

OpenStudy (anonymous):

\[f \circ f^{-1}(x)=x\] A function and it's inverse cancel each other out, leaving you with the independent variable x

OpenStudy (jhannybean):

I like the way @UnkleRhaukus proved it and then found the value.

OpenStudy (anonymous):

mmhm, it has a beauty of it's own

OpenStudy (anonymous):

yes it does